OpenStudy (anonymous):

how fast should a man of 70 kg run so that his kinetic energy is 875J ?

OpenStudy (anonymous):

@ganeshie8 help please !

OpenStudy (anonymous):

\[K.E = \frac{ 1 }{ 2 }* m*v^2\]

OpenStudy (anonymous):

875 = 1/2 * 70 * v^2 v^2 = 25 v = 5m/s

OpenStudy (anonymous):

@Yahoo! ty :)

OpenStudy (anonymous):

Welcome..)